George walks 5.0 km south, 10 km east, 8 km north, and 6.0 km west. Determine the displacement. Both magnitude and direction
1
Expert's answer
2020-10-23T09:09:55-0400

Considering coordinate system where east is directed towards x axis, then

Net displacement = 10i - 6i + 8j - 5j

= 4i + 3j

Magnitude of net displacement:

"=\sqrt{4^2+3^2}"

"=\sqrt{25}"

= 5 km

Direction:

"=\frac{4}{5}i+\frac{3}{5}j"